Calculate the concentration of hydroxide
ions in 1.25M HCl
1
Expert's answer
2019-12-03T08:24:42-0500

Calculate the concentration of hydroxide ions in 1.25M HCl

As the molar concentration of HCl equals 1.25M, 1.25 mol of HCl are present in 1L of the solution. As a result, 1.25 mol of hydroxide ions are present in the same volume.

From here, mass of hydroxide ions present in the solution equals:

m(H+) = n / Mr,

where n – number of moles, Mr – molecular weight

m(H+) = n / Mr = 1.25 mol / 1 g/mol = 1.25 g

As a result, the concentration of hydrxide ions is:

c(H+) = m / V= 1.25 g / 1L = 1.25 g/L

Answer: 1.25 g/L